HIGH RESOLUTION SPECTRUM FOR THE A^{1}Σ_{u}^{+}–X^{1}Σ_{g}^{+} SYSTEM OF LITHIUM ISOTOPOMERS

\underline{XUEJUN WANG}, J. YANG, J.B. QI and A.M. LYYRA, Department of Physics, Temple University, Philadelphia, PA 19122; ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

On combining our over 40 vibrational bands for the A^{1}Σ_{u}^{+}–X^{1}Σ_{g}^{+} system of ^{6}Li_{2} and ^{6,7}Li_{2} sub-Doppler polarization spectrum with literature data from the Fourier transform spectrum of this system for ^{7}Li_{2} ,^{a} we have determined a new set of Le Roy - type molecular constants,^{b} which include mass-reduced Dunham constants and Born-Oppenheimer and JWKB breakdown (BOB) correction terms, describing these two electronic states of all three isotopomers. The preliminary results indicate that the BOB corrections are not significant for the X^{1}Σ_{g}^{+} state, but are significant for A^{1}Σ_{u}^{+} state.

aLinh D. Le and R.A. Bernheim, “Atlas of the Absorption Spectrum of ^{6}Li_{2} and ^{7}Li_{2} .”

bR.J. Le Roy, paper FB09 at the “52^{nd}Ohio State University International Symposium on Molecular Spectroscopy”, Columbus, Ohio (1997); manuscript in preparation.